Activity 2: Maximum Error E



Activity 2: Maximum Error E



Assuming that the samples come from a normally distributed population, find



the maximum error E given the following:



Given



Solution and Answer



1.



n = 10, X= 28, s = 4.0,



90% confidence



2. n = 16, X= 50, s = 4.2,



95% confidence



3. n = 25, X= 92.8, s = 2.6,



99% confidence



Assuming that the samples come from a normally distributed population, find




the maximum error E given the following:




Given




Solution and Answer




1.




n = 10, X= 28, s = 4.0,




90% confidence




2. n = 16, X= 50, s = 4.2,




95% confidence




3. n = 25, X= 92.8, s = 2.6,




99% confidence

1
Expert's answer
2022-06-13T23:56:46-0400

1.

A 90% confidence interval, "df=n-1=9" degrees of freedom has a

t-score (a critical value) "t_c=1.833113."

"ME=t_c\times\dfrac{s}{\sqrt{n}}""=1.833113\times\dfrac{4.0}{\sqrt{10}}=2.3187"



2.

A 95% confidence interval, "df=n-1=24" degrees of freedom has a

t-score (a critical value) "t_c=1.833113."

"ME=t_c\times\dfrac{s}{\sqrt{n}}""=2.131449\times\dfrac{4.2}{\sqrt{16}}=2.2380"



3.

A 99% confidence interval, "df=n-1=24" degrees of freedom has a

t-score (a critical value) "t_c= 2.79694."

"ME=t_c\times\dfrac{s}{\sqrt{n}}""= 2.79694\times\dfrac{2.6}{\sqrt{25}}=1.4544"